A new firm is developing its business plan. It will require $615,000 of assets, and it projects $450,000 of sales and $355,000 of operating costs for the first year. Management is reasonably sure of these numbers because of contracts with its customers and suppliers. It can borrow at a rate of 7.5%, but the bank requires it to have a TIE of at least 4.0, and if the TIE falls below this level the bank will call in the loan and the firm will go bankrupt. What is the maximum debt ratio (measured as debt/assets) the firm can use? (Hint: Find the maximum dollars of interest, then the debt that produces that interest, and then the related debt ratio.)
Answer #1

TIE = EBIT / Interest

EBIT = 450,000 - 355,000 = $95,000

4 = $95,000 / Interest

Interest = $95,000 / 4 = $23,750. This is the most interest tha the company can afford and ot go over the TIE of 4.

Since Loan is at 7.5%, amount of debt will be:

$23,750 / .075 = $316.666.70

Debt Ratio = $316,666.70 / 615,000 = 0.5149